OpenStudy (jenniferjuice):

What are the x− and y−intercepts of the graph of y = x2 + 3x − 10? x−intercepts: (−2, 0) and (5, 0) y−intercept: (0, 10) x−intercepts: (−2, 0) and (−5, 0) y−intercept: (0, 10) x−intercepts: (−2, 0) and (5, 0) y−intercept: (0, −10) x−intercepts: (2, 0) and (−5, 0) y−intercept: (0, −10)

OpenStudy (jenniferjuice):

@Abhisar @ikram002p @jim_thompson5910 @texaschic101 @zepdrix

jimthompson5910 (jim_thompson5910):

to find the y-intercept, plug in x = 0 and evaluate

jimthompson5910 (jim_thompson5910):

what do you get for the y-intercept?

OpenStudy (jenniferjuice):

i plug in x=0 in how?

OpenStudy (jenniferjuice):

okay dumb question i got it its y= -10

jimthompson5910 (jim_thompson5910):

so the y-intercept is the point (0,-10)

jimthompson5910 (jim_thompson5910):

to find the x-intercept(s), you plug in y = 0 and solve for x

jimthompson5910 (jim_thompson5910):

x^2 + 3x - 10 = y x^2 + 3x - 10 = 0 x = ?? or x = ??

jimthompson5910 (jim_thompson5910):

were you able to solve it or no?

OpenStudy (jenniferjuice):

erm no

jimthompson5910 (jim_thompson5910):

we must factor x^2 + 3x - 10 find two numbers that multiply to -10 and add to 3 the numbers are 5 and -2 so x^2 + 3x - 10 = (x+5)(x-2)

jimthompson5910 (jim_thompson5910):

x^2 + 3x - 10 = 0 (x+5)(x-2) = 0 x = ?? or x = ??

OpenStudy (jenniferjuice):

5 and -2?

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

x+5 = 0 x = ??

OpenStudy (jenniferjuice):

-3? idkk

jimthompson5910 (jim_thompson5910):

x+5 = 0 x+5-5 = 0-5 x = -5

jimthompson5910 (jim_thompson5910):

that's one x-intercept what is the other?

OpenStudy (jenniferjuice):

-2?

jimthompson5910 (jim_thompson5910):

x-2 = 0 x = ?

OpenStudy (jenniferjuice):

+2?

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

so the x-intercepts are (2,0) and (-5,0) the y-intercept is (0,-10) all of this is shown in the graph (see attached)

jimthompson5910 (jim_thompson5910):

the graph was created using GeoGebra http://www.geogebra.org/cms/en/